OpenStudy (anonymous):

Find the limit Lim as x approaches infinity (Square root x^2+7)

OpenStudy (anonymous):

\[\lim_{x\to \infty}\sqrt{x^2+7}\]?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

that would be infinity wouldn't it?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

just gets bigger and bigger and bigger and ...

OpenStudy (anonymous):

No, my assignment wants numbers.

OpenStudy (anonymous):

ok make it 12,000.000

OpenStudy (anonymous):

Make it a billion xD

OpenStudy (anonymous):

@iambatman is more generous than i am`

ganeshie8 (ganeshie8):

most likely some mistake in the expression, take a screenshot and attach if psble

OpenStudy (anonymous):

It won't let me post an attachment

OpenStudy (anonymous):

You can always take a screenshot and use sites like puush and stuff which give a direct link to the image.

OpenStudy (anonymous):

Its the middle equation http://grab.by/BDGA

ganeshie8 (ganeshie8):

try entering `\infty` or `infinity`

OpenStudy (anonymous):

It worked thank you!!

OpenStudy (anonymous):

It seems you got the next problem wrong to, there is an easy way to do those problems, you see which ever in the numerator and denominator has the higher power, that's the limit (integer) if that makes sense, so it would be -1/9.